# Chromatic numbers, Buchstaber numbers and chordality of Bier spheres

### v1

@@ 1. Introduction
This paper is concerned with proving general formulae for the chromatic and Buchstaber numbers of Bier spheres as well as studying their chordality. Our interest in these particular combinatorial invariants and properties is motivated by their relations with each other as well as their applications in topology, algebra and combinatorics.
In the unpublished note [7], any simplicial complex $K$ on $[m]:=\{1,2,\ldots,m\}$ with $m\geq 2$, different from the whole simplex $\Delta_{[m]}$ on $[m]$, was associated with a simplicial complex $\mathrm{Bier}(K)$ defined as the deleted join of $K$ and its Alexander dual $K^{\vee}$. Moreover, it was shown in [7] that $\mathrm{Bier}(K)$ is an $(m-2)$-dimensional PL-sphere with a number of vertices varying between $m$ and $2m$. Since that time this beautiful and simple construction has been studied intensively and found important applications in topological combinatorics [24], polytope theory [19], optimization theory [18], game theory [28], combinatorial commutative algebra [17, 25] and toric topology [21, 22]. New proofs of the fact that it yields a sphere were also obtained, see [23, 24, 25].
Combinatorial and geometrical methods were used to classify all Bier spheres of dimension less than three [21, 22], all flag Bier and Murai spheres [17, 22] and to show that Bier spheres and their generalizations are shellable and edge decomposable [8, 25] as well as that asymptotically the vast majority of them are non-polytopal [24], although no particular examples of a non-polytopal Bier sphere have been found so far. Furthermore, methods of homological algebra allowed us to compute certain multigraded Betti numbers of Bier spheres and their generalizations [17, 25]. Finally, the application of topological combinatorics methods showed that all Bier spheres are starshaped [18], and methods of toric topology yielded the classification of all minimally non-Golod Bier spheres [22].
@@
The first combinatorial invariant of a Bier sphere studied in this paper is the chromatic number discussed in Section 3. By definition, this is the least number of colors needed for a proper coloring of the vertex set of a simplicial complex; that is, vertices of each simplex are colored in pairwise different colors. The chromatic number is one of the most important classical combinatorial invariants of a simple graph as well as of a simple polytope. By definition, the chromatic number $\chi(P)$ of an $n$-dimensional convex simple polytope $P$ equals that of its nerve complex $\partial P^{*}$ and hence $m\geq\chi(P)\geq n$, where $m$ is the number of facets of $P$. A combinatorial description of $n$-colorable simple $n$-polytopes was obtained in [20], while that in the case of $(n+1)$-colorable simple $n$-polytopes remains unknown.
Recall that a triangulated sphere is called polytopal if it is simplicially isomorphic to the boundary of a convex simplicial polytope. We prove that for any simplicial complex $K\neq\Delta_{[m]}$ on $[m]$ with $m\geq 2$ the chromatic number $\chi(\mathrm{Bier}(K))$ equals either $m-1$, or $m$. More precisely, the following statement holds, see Theorem 3.16.
@@
    $\chi(\mathrm{Bier}(K))=m-1$ if and only if $\mathrm{Bier}(K)$ is polytopal and one of the next two cases holds:
    • $P_{K}=I^{m-1}$;
    • $P_{K}=I^{m-3}\times P_{6}$.
Here, we denote by $P_{K}$ a simple polytope such that for a polytopal Bier sphere $\mathrm{Bier}(K)$ one has a combinatorial equivalence $\mathrm{Bier}(K)=\partial P_{K}^{*}$. We also use the notations $I^{n},n\geq 0$ for the standard $n$-dimensional cube and $P_{m},m\geq 3$ for an $m$-gon. Moreover, Lemma 3.15 shows that for $m\geq 3$, the chromatic number $\chi(\mathrm{Bier}(K))=m-1$ if and only if $K$ is obtained starting from one of the following three simplicial complexes on the set $[3]$: either (i) two points forming a segment complex, or (ii) three points forming a chain complex, or (iii) three disjoint points complex by iterative application of taking a cone and taking Alexander dual operations.
@@
The second combinatorial invariant of a Bier sphere studied in this paper is the Buchstaber number discussed in Section 4. By definition, the (complex) Buchstaber number of a simplicial complex $K$ with $m$ vertices is the maximal integer $r$ such that there exists a characteristic map $\Lambda\colon[m]\to\mathbb{Z}^{m-r}$; that is, each simplex of $K$ is mapped to a part of a lattice basis of $\mathbb{Z}^{m-r}$. We denote it by $s(K)$. Replacing $\mathbb{Z}$ by $\mathbb{Z}_{p}$ for a prime $p\in\mathbb{N}$ and the integer lattice by a $\mathbb{Z}_{p}$-vector space above, we get a definition of the mod $p$ Buchstaber number of $K$. We denote it by $s_{p}(K)$. The theory of complex and real Buchstaber numbers has been developed in the framework of toric topology and combinatorial commutative algebra, see [1, 2, 12, 13, 4, 5].
@@
These invariants play a crucial role in toric topology [9, 10] as they reflect important topological properties of polyhedral products [3]. Namely, the complex Buchstaber number of $K$ equals the maximal rank of a toric subgroup $H$ in the compact torus $(\mathbb{S}^{1})^{m}\subset({\mathbb{D}}^{2})^{m}$ such that the restricted coordinatewise action of $H$ on the moment-angle-complex $\mathcal{Z}_{K}\subseteq({\mathbb{D}}^{2})^{m}$ is free. Similarly one can interpret the mod $p$ Buchstaber numbers, see [6].
@@
It turns out that chromatic and Buchstaber numbers of an $(n-1)$-dimensional simplicial complex $K$ with $m$ vertices are closely related to each other by the following inequalities:
$m-\chi(K)\leq s(K)\leq s_{p}(K)\leq m-n,$
for any prime $p\in\mathbb{N}$, see [6, Corollary 5.8]. The real ($p=2$) and complex Buchstaber invariants of $\mathrm{Bier}(K)$ were computed in [21, Theorem 3.4] in the assumption that all the ghost vertices of $K$ are taken into account. In  Theorem 4.2 we get rid of that restriction proving the following result, which provides a general formula for all the Buchstaber numbers of $\mathrm{Bier}(K)$ only in terms of the $f$-vector of $K$.
@@
$s(\mathrm{Bier}(K))=s_{p}(\mathrm{Bier}(K))=f_{0}(K)-f_{m-2}(K)+1,\text{ for any prime }p\in\mathbb{N}.$
Since $\dim(\mathrm{Bier}(K))=m-2$ for any simplicial complex $K\neq\Delta_{[m]}$ on $[m]$ with $m\geq 2$, polytopal Bier spheres give rise to $(m-1)$- and $m$-colorable $(m-1)$-dimensional simple polytopes, which admit characteristic maps and therefore quasitoric manifolds and small covers due to [11].
Finally, chordality of Bier spheres is discussed in Section 4. Recall that a simple graph is chordal if it contains no induced subgraphs being cycles of length greater than three. We call a simplicial complex chordal if its 1-skeleton is a chordal graph. It turns out that this combinatorial property acquires algebraic and topological interpretations in the framework of toric topology.
@@
In Theorem 4.7 we classify all chordal Bier spheres by proving the following result.
@@
Let $K\neq\Delta_{[m]}$ be a simplicial complex on $[m]$ with $m\geq 2$. Then $\mathrm{Bier}(K)$ is chordal if and only if $\mathrm{Bier}(K)$ is polytopal and one of the next three cases holds:
@@
  • $P_{K}=\mathrm{vc}^{k}(\Delta^{m-1})$, where $0\leq k\leq m$ and $m\geq 4$.
Here, by $\mathrm{vc}^{k}(\Delta^{m-1})$ we mean a truncation polytope; that is, a simple polytope obtained from $\Delta^{m-1}$ by a sequence of $k$ vertex cuts by hyperplanes in general position. Moreover, as Lemma 4.6 asserts, in case $m\geq 4$, chordality of $\mathrm{Bier}(K)$ is equivalent to either $K$ or $K^{\vee}$ having no edges. This yields a canonical realization for the corresponding Bier sphere as a boundary of a stacked polytope; that is, one gets $\mathrm{Bier}(K)$ in that case as a boundary of a simplicial $(m-1)$-dimensional polytope obtained from the simplex $\Delta^{m-1}$ by performing stellar subdivisions (glueing pyramids) over some of its facets.
@@ 2. Basic definitions and constructions
In this section, we introduce our notations and recall some necessary definitions, examples and results related to the Bier sphere construction as well as to the notion of a chromatic number for simple graphs and, more generally, for simplicial complexes.
@@
Set $[m]:=\{1,2,\ldots,m\}$ for an integer $m\geq 1$. We say that a non-empty subset $K\subseteq 2^{[m]}$ is an abstract simplicial complex on $[m]$ if the following condition holds:
@@
One of the most important combinatorial invariants (i.e. characteristics invariant under a face lattice isomorphism) of a simplicial complex is its face vector. Namely, the $f$-vector of a simplicial complex $K$ on $[m]$ is a tuple $f(K):=(f_{-1}(K),\ldots,f_{m-1}(K))$, where $f_{i}(K)$ denotes the number of faces of dimension $i$ in $K$. Note that $f_{-1}(K)=1$ for any simplicial complex $K$, since $\varnothing\in K$ is the only simplex in $K$ of dimension $-1$, and $f_{i}(K)=0$ if $i>\dim(K)$.
@@
Let $m\geq 3$ and $P_{m}\subset{\mathbb{R}}^{2}$ be an $m$-gon. Then its dual polytope $P_{m}^{*}$ is isomorphic to $P_{m}$. Hence, $P_{m}$ is a simple and a simplicial polytope at the same time. The nerve complex of $P_{m}$ equals its boundary $\partial P_{m}$, which is the $m$-cycle denoted in what follows by $Z_{m}$.
@@
the full subcomplex in $K$ on the set $I$. In particular, if $i\in[m]$, then the deletion of $i$ from $K$ is the full subcomplex $\operatorname{Del}_{K}(i)$ in $K$ on the set $[m]\setminus\{i\}$. If $I\in K$ we call the simplicial complex
@@
$V(\mathrm{sk}^{0}(K))=V(K)=V(\mathrm{sk}^{1}(K))\text{ and }\mathrm{sk}^{0}(K)=V(K)\sqcup\{\varnothing\},$
@@
  • $\Sigma_{v,w}(K):=\partial\Delta_{\{v,w\}}\ast K$, the suspension over $K$ with vertex pair $\{v,w\}$.
@@
Now we are going to introduce the main object of study in this paper, a Bier sphere. In the definition below, we use the following notations. Let $[m]:=\{1,2,\ldots,m\}$ and $[m^{\prime}]:=\{1^{\prime},2^{\prime},\ldots,m^{\prime}\}$ be two disjoint ordered sets with the map $\phi\colon i\mapsto i^{\prime},1\leq i\leq m$ being an order preserving bijection between them. Denote by $I^{\prime}$ the image $\phi(I)$ of a subset $I\subseteq[m]$.
@@
  • the Alexander dual of $K$ is a simplicial complex $K^{\vee}$ on $[m^{\prime}]$ such that
@@
  • the Bier sphere of $K$ is a simplicial complex $\mathrm{Bier}(K)$ on $[m]\sqcup[m^{\prime}]$ such that
@@
It was shown in [7] that $\mathrm{Bier}(K)$ defined above is a PL-sphere of dimension $(m-2)$ with the number of vertices varying between $m$ and $2m$. It follows from the definition that the Bier sphere construction is symmetric: $\mathrm{Bier}(K)=\mathrm{Bier}(K^{\vee})$. In Figure 1 one can find the geometric realizations of all possible 1-dimensional Bier spheres, see also [22, Example 2.8].
@@
Thus, we see that both $\mathrm{Bier}(K_{1})$ and $\mathrm{Bier}(K_{2})$ are $(m-3)$-times iterated suspensions over the 4-cycle $Z_{4}$, or, in other words, both are isomorphic to the boundary of the $(m-1)$-dimensional cross-polytope $(I^{m-1})^{*}$.
Taking into account Proposition 2.9 and Example 2.11, our first goal is to study the chromatic numbers of Bier spheres. When $m=2$, the only Bier sphere has a geometric realization as the set of two disjoint points, it has chromatic number $1$. Therefore, in what follows, if otherwise not stated explicitly, we always assume that a simplicial complex on $[m]$ under consideration is different from $\Delta_{[m]}$ and that $m$ is always greater or equal to $3$.
@@
Figure 2. The complexes $G_{4}$ and $\Gamma_{6}$ are self-dual and $G_{4}$ has a ghost vertex. The indices in the notation of a complex indicate the number of vertices in the corresponding Bier sphere.
The chromatic number of Bier spheres $\mathrm{Bier}(\Gamma_{4})$, $\mathrm{Bier}(G_{4})$, and $\mathrm{Bier}(\Gamma_{6})$ is 2, since it is $Z_{4}$ in the first two cases and $Z_{6}$ in the last case. The chromatic number of $\mathrm{Bier}(\Gamma_{3})$ and $\mathrm{Bier}(\Gamma_{5})$ is 3, since they are isomorphic to $Z_{3}$ and $Z_{5}$, respectively. This motivates us to compute the chromatic number of an arbitrary Bier sphere: it is done in the next section.

### v2

@@ 1. Introduction
This paper is a part of a program initiated in [27], with the goal of constructing smooth moment-angle manifolds and toric manifolds over Bier spheres and applying them to solve a few important problems in algebraic and equivariant topology. Moment-angle manifolds and (quasi)toric manifolds were introduced in [13] and have become the main objects of study in toric topology [10]. It is a modern area at the crossroads of algebraic and equivariant topology, where combinatorial properties of an abstract simplicial complex $K$ on $m]:=\{1,2,\ldots,m\}$ are studied in their relationship with the topological properties of the corresponding moment-angle complex $\mathcal{Z}_{K}\subset\mathbb{C}^{m}$, its real counterpart $\mathcal{R}_{K}\subset\mathbb{R}^{m}$, as well as their orbit spaces with respect to a freely acting subtorus $H$ in ${\mathbb{T}}^{m}$ and $\mathbb{Z}_{2}^{m}$, respectively. The latter class of partial quotients includes (quasi)toric manifolds and their real counterparts as the subclass, where the toric subgroup $H$ has a maximal possible rank.
Quasitoric manifolds found numerous applications in algebraic and equivariant topology, with probably the most remarkable one being that they yield geometric representatives in unitary [11] and special unitary [12] bordism ring classes. However, quasitoric manifolds are not necessarily toric and therefore do not solve the Hirzebruch problem in algebraic topology: find a nonsingular connected algebraic variety representing a given unitary bordism class. One of our goals in the above mentioned program is to construct nonsingular complete toric varieties (toric manifolds) over Bier spheres and describe all the unitary bordism classes in which there exists such a toric manifold. Other goals will also be described in this section of the paper.
Recall that in the unpublished note [7], any (abstract) simplicial complex $K$ on $[m]$ with $m\geq 2$, different from the whole simplex $\Delta_{[m]}$ on $[m]$, was associated with a simplicial complex $\mathrm{Bier}(K)$ defined as the deleted join of $K$ and its Alexander dual $K^{\vee}$. Moreover, it was shown in [7] that $\mathrm{Bier}(K)$ is an (m2)(m-2)-dimensional PL-sphere with a number of vertices varying between $m$ and $2m$. Since that time, this beautiful and simple construction has been studied intensively and found important applications in topological combinatorics [30], polytope theory [24], optimization theory [23], game theory [34], combinatorial commutative algebra [21, 31] and toric topology [27, 28]. New proofs of the fact that it yields a sphere were also obtained, see [29, 30, 31].
Combinatorial and geometrical methods were used to classify all Bier spheres of dimension less than three [27, 28], all flag Bier and Murai spheres [21, 28] and to show that Bier spheres and their generalizations are shellable and edge decomposable [8, 31] as well as that asymptotically the vast majority of them are non-polytopal [30], although no particular examples of a non-polytopal Bier sphere have been found so far. Furthermore, methods of homological algebra made it possible to compute certain multigraded Betti numbers of Bier spheres and their generalizations [21, 31]. Finally, the application of topological combinatorics methods showed that all Bier spheres are starshaped [23], and methods of toric topology yielded the classification of all minimally non-Golod Bier spheres [28].
@@
The first combinatorial invariant of a Bier sphere studied in this paper is the chromatic number discussed in Section 3. By definition, this is the least number of colors needed for a proper coloring of the vertex set of a simplicial complex; that is, the vertices of each simplex are colored in pairwise different colors. The chromatic number is one of the most important classical combinatorial invariants of a simple graph as well as of a simple polytope. By definition, the chromatic number $\chi(P)$ of an $n$-dimensional convex simple polytope $P$ equals that of its nerve complex $\partial P^{*}$ and hence $m\geq\chi(P)\geq n$, where $m$ is the number of facets of $P$. A combinatorial description of $n$-colorable simple $n$-polytopes was obtained in [25]. Computation of chromatic numbers for families of simple polytopes is a classical open problem in graph theory, polytope theory and applications, see [5].
Recall that a triangulated sphere is called polytopal if it is simplicially isomorphic to the boundary of a convex simplicial polytope. We prove that for any simplicial complex $K\neq\Delta_{[m]}$ on $[m]$ with $m\geq 2$ the chromatic number $\chi(\mathrm{Bier}(K))$ equals either m1m-1, or $m$. More precisely, the following statement holds, see Theorem 3.16.
@@
    $\chi(\mathrm{Bier}(K))=m-1$ if and only if $\mathrm{Bier}(K)$ is polytopal and the next equivalent conditions hold:
    • $K\in\{C^{m-3}(\Gamma_{4})$, $C^{m-3}(\Gamma_{4}^{\vee})$, $C^{m-3}(G_{4})$, $C^{m-3}(\Gamma_{6})\}$;
    • $P_{K}\in\{I^{m-1},I^{m-3}\times P_{6}\}$.
Here, we use the notation $C^{\ell}(\mathcal{K})$ for the $\ell$ times iterated cone over a simplicial complex $\mathcal{K}$ and the corresponding 4 graphs are described in Figure 2. We also denote by $P_{K}$ a simple polytope such that for a polytopal Bier sphere $\mathrm{Bier}(K)$ one has a combinatorial equivalence $\mathrm{Bier}(K)=\partial P_{K}^{*}$ and use the notations $I^{n}$, $n\geq 0$ for the standard $n$-dimensional cube and $P_{m}$, $m\geq 3$ for an $m$-gon. A combinatorial classification of simple $n$-polytopes $P$ with $\chi(P)=n+1$ remains unknown in polytope theory. Our theorem implies that every Bier polytope $P_{K}$ with $K$ different from each of the complexes described explicitly in Theorem 3.16 belongs to this class and, in view of the result of [25], it has a 2-dimensional face with an odd number of vertices.
@@
The second combinatorial invariant of a Bier sphere studied in this paper is the Buchstaber number discussed in Section 4. By definition, the (complex) Buchstaber number of a simplicial complex $K$ with $m$ vertices is the maximal integer $r$ such that there exists a characteristic map $\Lambda\colon[m]\to\mathbb{Z}^{m-r}$; that is, each simplex of $K$ is mapped to a part of a lattice basis of $\mathbb{Z}^{m-r}$. We denote it by $s(K)$. Replacing $\mathbb{Z}$ by $\mathbb{Z}_{p}$ for a prime $p\in\mathbb{N}$ and the integer lattice by a $\mathbb{Z}_{p}$-vector space above, we get a definition of the mod $p$ Buchstaber number of $K$. We denote it by $s_{p}(K)$. The theory of complex and real Buchstaber numbers has been developed in the framework of toric topology and combinatorial commutative algebra, see [1, 2, 15, 17, 4].
@@
These invariants play a crucial role in toric topology [9, 10] as they reflect important topological properties of polyhedral products [3]. Namely, the complex Buchstaber number of $K$ equals the maximal rank of a toric subgroup $H$ in the compact torus ${\mathbb{T}}^{m}\subset({\mathbb{D}}^{2})^{m}$ such that the restricted coordinatewise action of $H$ on the moment-angle-complex $\mathcal{Z}_{K}\subseteq({\mathbb{D}}^{2})^{m}$ is free. Similarly, one can interpret the mod $p$ Buchstaber numbers, see [6].
@@
It turns out that chromatic and Buchstaber numbers of an (n1)(n-1)-dimensional simplicial complex $K$ with $m$ vertices are closely related to each other by the following inequalities:
$m-\chi(K)\leq s(K)\leq s_{p}(K)\leq m-n\,\text{ for any prime }p\in\mathbb{N},$
which we discuss and use in Section 4. The case of the maximal Buchstaber invariant s(K)=mns(K)=m-n and a starshaped sphere $K$ is of particular importance in toric topology; then the $(m+n)$-dimensional cellular space $\mathcal{Z}_{K}$ acquires an equivariant smooth structure and the above mentioned orbit space construction yields a stably complex $2n$-dimensional closed, orientable manifold $M(K,\Lambda)=\mathcal{Z}_{K}/H$, which generalizes the notion of a quasitoric manifold corresponding to a polytopal sphere $K$, see [10]. Classification of these manifolds and, in particular, a combinatorial description of all starshaped spheres with a maximal Buchstaber number remains an open problem of equivariant topology.
In Section 4 we show that all Bier spheres belong to this class of starshaped spheres, which gives rise to a new wide class of quasitoric manifolds. Indeed, due to [15], $\chi(K)\in\{n,n+1\}$ for a polytopal sphere $K$ implies that s(K)=mns(K)=m-n; then Theorem 3.16 (in the polytopal case) and our construction of a canonical characteristic map (in the general case) yield the following general formula, see Theorem 4.2, for all the Buchstaber numbers of $\mathrm{Bier}(K)$ in terms of the $f$-vector of $K$ only.
@@
In our subsequent publications, we are going to use the previous result in order to show that our canonical characteristic map $\Lambda_{K}$ for a Bier sphere $\mathrm{Bier}(K)$ determines the rays of a complete regular fan, whose underlying simplicial complex is isomorphic to $\mathrm{Bier}(K)$. Therefore, the corresponding manifold $M(\mathrm{Bier}(K),\Lambda_{K})$ acquires a structure of a toric manifold. Taking $\mathrm{Bier}(K)$ to be a non-polytopal Bier sphere, this immediately solves another problem of equivariant topology: to show the existence of infinitely many toric, but not quasitoric manifolds.
Finally, the chordality of Bier spheres is discussed in Section 4. Recall that a simple graph is chordal if it contains no induced subgraphs being cycles of length greater than three. We call a simplicial complex chordal if its 1-skeleton is a chordal graph. It turns out that this combinatorial property acquires the following algebraic and topological interpretations in the framework of toric topology.
We are interested here in the following open problem of algebraic topology: classify all closed connected manifolds $M$ with $\mathrm{cat}(M)=2$, where $\mathrm{cat}(X)$ denotes the Lusternik-Schnirelmann category of a topological space $X$. It is well-known that $\mathrm{cat}(X)=1$ if and only if $X$ is a co-H-space, hence $\mathrm{cat}(\mathcal{Z}_{K})=1$ for a starshaped sphere $K$ if and only if $K$ is a boundary of a simplex. Thus, our next goal is to describe the class of moment-angle manifolds $\mathcal{Z}_{K}$ with $\mathrm{cat}(\mathcal{Z}_{K})=2$ in combinatorial and algebraic terms. By the result of [14], a closed connected $n$-manifold $M$ with $n>2$ has $\mathrm{cat}(M)=2$ only if $\pi_{1}(M)$ is a free group. Furthermore, it was shown in [33] that the fundamental group $\pi_{1}(\mathcal{R}_{K})$ of a real moment-angle-complex $\mathcal{R}_{K}$ is isomorphic to the commutator subgroup in the right-angled Coxeter group of $\mathrm{sk}^{1}(K)$ and the latter is a free group if and only if $K$ is chordal. Therefore, given a starshaped sphere $K$ of dimension n1>1n-1>1, if the real moment-angle manifold $\mathcal{R}_{K}$ of dimension $n$ has $\mathrm{cat}(\mathcal{R}_{K})=2$, then $K$ is chordal.
[…]
Conjecture 1.3.
Let $K$ be an (n1)(n-1)-dimensional chordal starshaped sphere with $n>2$. Then the next conditions are equivalent:
  1. (1)
    $\mathrm{cat}(\mathcal{Z}_{K})=2$;
  2. (2)
    $\mathrm{cat}(\mathcal{R}_{K})=2$;
  3. (3)
    $\mathcal{Z}_{K}$ is homeomorphic to a connected sum of sphere products with two spheres in each product;
  4. (4)
    $\mathcal{R}_{K}$ is homeomorphic to a connected sum of sphere products with two spheres in each product;
  5. (5)
[…]
In this paper, we are going to make a first step towards the proof of the above conjecture in the class of moment-angle manifolds over Bier spheres. Namely, in Theorem 4.7, we classify all chordal and all stacked Bier spheres by proving the following result.
@@
Let $K\neq\Delta_{[m]}$ be a simplicial complex on $[m]$ with $m\geq 2$. Then the following three statements are equivalent:
  • (a)
    $\mathrm{Bier}(K)$ is a chordal complex;
  • (b)
    $\mathrm{Bier}(K)$ is a polytopal sphere and one of the next three cases holds:
    1. (1)
@@
    3. (3)
      $P_{K}\in\mathrm{vc}^{k}(\Delta^{m-1})$, where $0\leq k\leq m$ and $m\geq 4$.
  • (c)
    One of the next three cases holds:
    1. (1)
      $m=2$;
    2. (2)
      $m=3$ and either $K$ or $K^{\vee}$ equals $\partial\Delta^{2}$;
    3. (3)
      $m\geq 4$ and either $K$ or $K^{\vee}$ has no edges.
If $m\leq 3$, then $\mathrm{Bier}(K)$ is a stacked sphere being either the boundary of a simplex, or the $k$-cycle with $k=4,5,$ or $6$. If $m\geq 4$, then $P_{K}$ is obtained by cutting off $k$ vertices of the simplex $\Delta^{m-1}$, where $k$ is the number of vertices of the corresponding complex without edges; in this case, (a), (b) and (c) are equivalent to:
  • (d)
    $\mathrm{Bier}(K)$ is a stacked sphere.
Here, by $\mathrm{vc}^{k}(\Delta^{m-1})$ we mean a set of (combinatorial) truncation polytopes; each of them can be obtained from $\Delta^{m-1}$ by a sequence of $k$ vertex cuts by hyperplanes in general position. It follows from [26, Theorem 8.5] that a simplicial (d1)(d-1)-sphere with $d\geq 3$ is a stacked sphere (i.e. is isomorphic to the nerve complex of a truncation polytope) if and only if it is chordal and has no minimal non-faces of cardinality $k$ for $2<k<d$. Obviously, in case $m\leq 3$, all Bier spheres are stacked, see [28, Example 2.8]. As Lemma 4.6 asserts, in case $m\geq 4$, the chordality of $\mathrm{Bier}(K)$ is equivalent to either $K$ or $K^{\vee}$ having no edges. This yields a canonical realization for the corresponding Bier sphere as a boundary of a stacked polytope; that is, one gets $\mathrm{Bier}(K)$ in that case as a boundary of a simplicial (m1)(m-1)-dimensional polytope obtained from the simplex $\Delta^{m-1}$ by performing stellar subdivisions (gluing pyramids) over some of its facets. Based on this classification, we are going to prove Conjecture 1.3 for all chordal Bier spheres in our subsequent publications.
@@ 2. Basic definitions and constructions
In this section, we introduce our notations and recall some necessary definitions, examples, and results related to the Bier sphere construction as well as to the notion of chromatic number for simple graphs and, more generally, for simplicial complexes.
@@
Set $m]:=\{1,2,\ldots,m\}$ for an integer $m\geq 1$. We say that a non-empty subset $K\subseteq 2^{[m]}$ is an abstract simplicial complex on $[m]$ if the following condition holds:
@@
One of the most important combinatorial invariants (i.e. characteristics invariant under a face lattice isomorphism) of a simplicial complex is its face vector. Namely, the $f$-vector of a simplicial complex $K$ on $[m]$ is a tuple $f(K):=(f_{-1}(K),\ldots,f_{m-1}(K))$, where $f_{i}(K)$ denotes the number of faces of dimension $i$ in $K$. Note that $f_{-1}(K)=1$ for any simplicial complex $K$, since $\varnothing\in K$ is the only simplex in $K$ of dimension 1-1, and $f_{i}(K)=0$ if $i>\dim(K)$.
@@
Let $m\geq 3$ and $P_{m}\subset{\mathbb{R}}^{2}$ be an $m$-gon. Then its dual polytope $P_{m}^{*}$ is isomorphic to $P_{m}$. Therefore, $P_{m}$ is a simple and a simplicial polytope at the same time. The nerve complex of $P_{m}$ equals its boundary $\partial P_{m}$, which is the $m$-cycle denoted in what follows by $Z_{m}$.
@@
the full subcomplex in $K$ on the set $I$. In particular, if $i\in[m]$, then the deletion of $i$ from $K$ is the full subcomplex $\operatorname{Del}_{K}(i)$ in $K$ on the set $m]\setminus\{i\}$. If $I\in K$ we call the simplicial complex
@@
[…]
@@
  • $\Sigma_{\{v,w\}}(K):=\partial\Delta_{\{v,w\}}\ast K$, the suspension over $K$ with vertex pair $\{v,w\}$.
@@
Now we are going to introduce the main object of study in this paper, a Bier sphere. In the definition below, we use the following notations. Let $m]:=\{1,2,\ldots,m\}$ and $m^{\prime}]:=\{1^{\prime},2^{\prime},\ldots,m^{\prime}\}$ be two disjoint ordered sets with the map $\phi\colon i\mapsto i^{\prime},1\leq i\leq m$ being an order preserving bijection between them. Denote by $I^{\prime}$ the image $\phi(I)$ of a subset $I\subseteq[m]$.
@@
  • the Alexander dual of $K$ is a simplicial complex $K^{\vee}$ on $m^{\prime}]$ such that
@@
  • the Bier sphere of $K$ is a simplicial complex $\mathrm{Bier}(K)$ on $m]\sqcup[m^{\prime}]$ such that
@@
It was shown in [7] that $\mathrm{Bier}(K)$ defined above is a PL-sphere of dimension (m2)(m-2) with the number of vertices varying between $m$ and $2m$. It follows from the definition that the Bier sphere construction is symmetric: $\mathrm{Bier}(K)=\mathrm{Bier}(K^{\vee})$. In Figure 1 one can find the geometric realizations of all possible 1-dimensional Bier spheres, see also [28, Example 2.8].
@@
Thus, we see that both $\mathrm{Bier}(K_{1})$ and $\mathrm{Bier}(K_{2})$ are (m3)(m-3)-times iterated suspensions over the 4-cycle $Z_{4}$, or, in other words, both are isomorphic to the boundary of the (m1)(m-1)-dimensional cross-polytope $(I^{m-1})^{*}$.
Taking into account Proposition 2.9 and Example 2.11, our first goal is to study the chromatic numbers of Bier spheres. When $m=2$, the only Bier sphere has a geometric realization as the set of two disjoint points, it has chromatic number $1$. Therefore, in what follows, if otherwise not stated explicitly, we always assume that a simplicial complex on $[m]$ under consideration is different from $\Delta_{[m]}$ and that $m$ is always greater than or equal to $3$.
@@
Figure 2. The complexes $G_{4}$ and $\Gamma_{6}$ are self-dual, $G_{4}$ has a ghost vertex. The indices in the notation of a complex indicate the number of vertices in the corresponding Bier sphere.
The chromatic number of Bier spheres $\mathrm{Bier}(\Gamma_{4})$, $\mathrm{Bier}(G_{4})$, and $\mathrm{Bier}(\Gamma_{6})$ is 2, since it is $Z_{4}$ in the first two cases and $Z_{6}$ in the last case. The chromatic number of Bier spheres $\mathrm{Bier}(\Gamma_{3})$ and $\mathrm{Bier}(\Gamma_{5})$ is 3, since they are isomorphic to $Z_{3}$ and $Z_{5}$, respectively. This motivates us to compute the chromatic number of an arbitrary Bier sphere: it is done in the next section.